OpenStudy (anonymous):

verify 1/csc^2(x)+1/sec^2(x)=1

OpenStudy (anonymous):

csc^2(x)=1/sin^2(x) sec^2(x)=1/cos^2(x) 1/csc^2(x)=sin^2(x) 1/sec^2(x)=cos^2(x) sin^2(x)+cos^2(x)=1 which is a trig identity